The marks of students in a mathematical contest is taken to follow N( 60,100 ). A team has 3 students. Students are rewarded if their measure for the team is 75 and above. Find the probability that a team is rewarded
1
Expert's answer
2020-12-21T18:33:39-0500

Let "X=" the mark of student in a mathematical contest: "X\sim N(\mu,\sigma^2/n)." Then

"Z=\dfrac{X-\mu}{\sigma/\sqrt{n}}\sim N(0,1)"

Given "\mu=60, \sigma^2=100, n=3"


"P(X\geq75)=1-P(X<75)"

"=1-P(Z<\dfrac{75-70}{10/\sqrt{3}})"

"\approx1-P(Z<0.866)"

"\approx1-0.806762=0.193238"



The probability that a team is rewarded is "0.193238."